equal amount of an Ideal diatomic gas is taken in two cylinders a and b at 300K. The Piston of A is free to move while that of B is it fixed. Same quantity of heat is given to both the cylinders. If the rise in temperature of gas is a is 30 k, the rise in temperature of gas in B is 
In this solution how to calculate Cp/Cv

Grade:11

1 Answers

Forum 1
one year ago
Hello Student

Solution:
Process in cylinder A is constant pressure process.
Process in cylinder B is constant volume process.
Heat supplied is :
Q=nCp​ΔT1​
Q=nCv​ΔT2​
Cp/Cv​​​=​ΔT2/ΔT1​​=γ
For a monoatomic gas is
γ=7/5
HenceT2​=7/5​×30=42K

Hope this solution is helpful.